# Variational Prior Replacement in Bayesian Inference and Inversion Xuebin Zhao1 and Andrew Curtis1

### v1

@@ 5 Conclusions
We develop a variational prior replacement (VPR) methodology designed to efficiently update prior information in Bayesian inference solutions. This approach involves replacing the existing prior information in a posterior distribution obtained from a previous inference process with a new prior distribution. The new posterior distribution is then found using variational inference. VPR eliminates the need to re-solve Bayesian inverse problems from scratch each time prior information changes. In a 2D full waveform inversion example, we prove the effectiveness of VPR for varying prior information, in which VPR provides a nearly identical posterior probability distribution compared to that obtained using the conventional prior specific inversion scheme. This similarity holds for individual posterior samples, as well as first- and second-order statistics. The key advantage of VPR lies in its computational efficiency: achieving the same results in a matter of minutes compared to two days required by the conventional approach. Additionally, we show that VPR can be used to investigate the impact of different prior distributions on Bayesian inference results. This methodology has significant potential for applications in more computationally demanding inverse problems, such as 3D Bayesian FWI, especially when multiple priors are available or need to be tested and discriminated for the same dataset.
@@ Acknowledgements.
- Kobyzev et al. (2019) Kobyzev, I., Prince, S., & Brubaker, M. A., 2019. Normalizing flows: An introduction and review of current methods, arXiv preprint arXiv:1908.09257.
@@
- Papamakarios et al. (2019) Papamakarios, G., Nalisnick, E., Rezende, D. J., Mohamed, S., & Lakshminarayanan, B., 2019. Normalizing flows for probabilistic modeling and inference, arXiv preprint arXiv:1912.02762.
@@
- Zhang & Curtis (2020) Zhang, X. & Curtis, A., 2020. Seismic tomography using variational inference methods, Journal of Geophysical Research: Solid Earth, 125(4), e2019JB018589.
@@
- Zhang et al. (2018) Zhang, X., Curtis, A., Galetti, E., & De Ridder, S., 2018. 3-d monte carlo surface wave tomography, Geophysical Journal International, 215(3), 1644–1658.

### v2

@@ 5 Conclusions
We develop a variational prior replacement (VPR) methodology designed to efficiently update prior information in Bayesian inference solutions. This approach involves replacing the existing prior information in a posterior distribution obtained from a previous inference process with a new prior distribution. The new posterior distribution is then found using variational inference. VPR eliminates the need to re-solve Bayesian inverse problems from scratch each time prior information changes. The results from a 2D full waveform inversion example support the effectiveness of VPR for varying prior information, in which VPR provides consistent statistics of the posterior probability distribution compared to those obtained using the conventional prior specific inversion scheme. This similarity holds for individual posterior samples, first- and second-order statistics, as well as simulated waveform data. The key advantage of VPR lies in its computational efficiency: achieving the same results in a matter of minutes compared to two days required by the conventional approach. Additionally, we show that VPR can be used to investigate the impact of different prior distributions on Bayesian inference results. This methodology has significant potential for applications in more computationally demanding inverse problems, such as 3D Bayesian FWI, especially when multiple priors are available or need to be tested and discriminated for the same dataset.
6 Data Availability
The code underlying this article will be shared on reasonable request to the corresponding author.
@@ Acknowledgements.
- Kobyzev et al. (2020) Kobyzev, I., Prince, S. J., & Brubaker, M. A., 2020. Normalizing flows: An introduction and review of current methods, IEEE transactions on pattern analysis and machine intelligence, 43(11), 3964–3979.
@@
- Papamakarios et al. (2021) Papamakarios, G., Nalisnick, E., Rezende, D. J., Mohamed, S., & Lakshminarayanan, B., 2021. Normalizing flows for probabilistic modeling and inference, Journal of Machine Learning Research, 22(57), 1–64.
@@
- Sun & Williamson (2024) Sun, Y. & Williamson, P., 2024. Invertible neural networks for uncertainty quantification in refraction tomography, The Leading Edge, 43(6), 358–366.
@@
- Zhang & Curtis (2020a) Zhang, X. & Curtis, A., 2020a. Seismic tomography using variational inference methods, Journal of Geophysical Research: Solid Earth, 125(4), e2019JB018589.
- Zhang & Curtis (2020b) Zhang, X. & Curtis, A., 2020b. Variational full-waveform inversion, Geophysical Journal International, 222(1), 406–411.
@@
Appendix A Verifying VPR using a Normal Initial prior Distribution
In this appendix, we present a second example to test the performance of the proposed variational prior replacement (VPR) methodology. We define a diagonal Gaussian distribution as the old prior pdf, with mean and standard deviation are calculated from the uniform prior distribution used in the main text. Figure 12a shows the inversion results obtained using this normal prior distribution. From top to bottom, each panel represents a random posterior sample, the mean velocity, standard deviation and the relative error maps of the posterior distribution, where the relative error is calculated as the difference between the mean and true velocity models divided by the standard deviation at each point. Given this inversion result, we perform VPR by replacing a smoothed version of the normal distribution (which is defined using exactly the same way as that expressed in equation 20) by the original normal prior distribution, and display the corresponding results in Figure 12c. Similarly to the main text, in Figure 12b we display prior specific inversion (PSI) results obtained using this same smoothed prior distribution. Since Figures 12b and 12c present highly consistent results, we conclude that VPR is accurate and effective in this example.